OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 24, 36, 54, 81, ... A. an=24(1.5)n-1 B. an=24(0.5)n-1 C. an=24(12)n-1 D. an=24(18)n-1

OpenStudy (anonymous):

A and formula would be an=24(1.5)^n-1

OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 11, 33, 99, 297, ... A. B. an=11(3)n-1 C. an=3(11)n-1 D.

OpenStudy (anonymous):

thius is answer a

OpenStudy (anonymous):

nd this is answer da

OpenStudy (anonymous):

an=11(3)^n-1